OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y-4x+10 y=1/4x +2/5 y=-4x - 6 y= -1/4x + 3/2

OpenStudy (anonymous):

is the first one y= -4x+10?

OpenStudy (anonymous):

Parallel lines have same slope. Which line has same slopwe with y=4x+7?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

okay so therefore it can be either 1 or 3

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

i say 3

OpenStudy (anonymous):

so you take each equation and you substitute the coordinates it into the two choices

OpenStudy (anonymous):

2=-4(-2)-6 2= 8 -6 2=2

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown. y = 1/3x y = -1/3x + 4 y = 3x + 20 y = -3x - 16

OpenStudy (anonymous):

perpendicular lines are negative reciprocals

OpenStudy (anonymous):

are you talking about the new problem i posted

OpenStudy (anonymous):

so its number 2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the line perpendicular to 3x would be -1/3x

OpenStudy (anonymous):

Consider the line that passes through the points (-5, 9) and (-2, 0). Below are two different equations. Identify the true statement concerning both equations. equation1: y-9=-1/3(x+5) equation2: y - 9 = -3(x + 5) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points

OpenStudy (anonymous):

@salgai

OpenStudy (anonymous):

sub in the coordinates into the equations

OpenStudy (anonymous):

and its pretty simple which statement is correct

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

let me check one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

aww how

OpenStudy (anonymous):

the slope of the line is -3 and only the first equation has a slope of -3

OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (anonymous):

Here are the steps 1) for all the possible answers, you will have to isolate the variable "y" 2) than you can check the slope for each, by checking the value of the constant before the x value (it has to be -3) 3) Now you have to substitute (7,-2) into the equations to see which ones fit. So, substitute 7 for x and -2 for y. 4) The correct answer is the one that balances.